# Dolphin: A Programmable Framework for Scalable Neurosymbolic Learning

### v1

@@ Abstract
Neurosymbolic learning has emerged as a promising paradigm to incorporate symbolic reasoning into deep learning models. However, existing frameworks are limited in scalability with respect to both the training data and the complexity of symbolic programs. We propose Dolphin, a framework to scale neurosymbolic learning at a fundamental level by mapping both forward chaining and backward gradient propagation in symbolic programs to vectorized computations. For this purpose, Dolphin introduces a set of abstractions and primitives built directly on top of a high-performance deep learning framework like PyTorch, effectively enabling symbolic programs to be written as PyTorch modules. It thereby enables neurosymbolic programs to be written in a language like Python that is familiar to developers and compile them to computation graphs that are amenable to end-to-end differentiation on GPUs. We evaluate Dolphin on a suite of 13 benchmarks across 5 neurosymbolic tasks that combine deep learning models for text, image, or video processing with symbolic programs that involve multi-hop reasoning, recursion, and even black-box functions like Python eval(). Dolphin only takes $0.33$%-$37.17$% of the time (and $2.77$% on average) to train these models on the largest input per task compared to baselines Scallop, ISED, and IndeCateR+, which time out on most of these inputs. Models written in Dolphin also achieve state-of-the-art accuracies even on the largest benchmarks.
@@ 1 Introduction
Deep learning has achieved remarkable success on a wide range of tasks such as image classification, speech recognition, and natural language processing. With the recent emergence of foundation models like GPT-4 and CLIP, deep learning models are now increasingly being used in more challenging tasks. While such models work well for prediction and generation tasks, they are limited in their ability to perform reasoning required for tasks involving structure, logic, and planning, where symbolic reasoning traditionally excels (Kambhampati et al., 2024). Neurosymbolic programming (Chaudhuri et al., 2021) has emerged as a promising paradigm to incorporate symbolic reasoning into deep learning models, providing the best of both worlds.
@@
Various frameworks have been developed to improve the programmability and accessibility of neurosymbolic applications (Manhaeve et al., 2018; Li et al., 2023; Solko-Breslin et al., 2024). These frameworks support complex symbolic reasoning features like recursion and black-box functions, implement efficient differentiable reasoning algorithms, and provide bindings for deep learning frameworks like PyTorch. Despite significant strides, however, these frameworks incur significant overhead when training neurosymbolic programs.
Consider a typical workflow of such a framework depicted in Figure 1(a). In this case, we have a supervised learning task with labeled data $(x,y)$, a neural network $M_{\theta}$ that processes input $x$, and a symbolic program $P_{\tt symbolic}$ that takes the network’s output $r$ and produces final output $y$. Existing frameworks, such as Scallop (Li et al., 2023), execute the neural model on GPU but use a separate CPU-based backend (implemented in Rust in Scallop’s case) for the symbolic program. Moreover, such systems suffer from slow inter-process data transfers between the two environments.
(a) Typical neurosymbolic framework (e.g. Scallop).
(b) Our neurosymbolic framework Dolphin.
Figure 1: Comparison of system architectures of neurosymbolic frameworks.
Together, these issues hinder the scalability of neurosymbolic learning with respect to problem complexity and data complexity. First, the symbolic computation engine must derive a set of all possible results and their associated probabilities in a manner that is differentiable with respect to the network’s parameters $\theta$. As the complexity of the symbolic program increases, the number of possible results and their associated weights also grows exponentially, leading to a combinatorial explosion in the number of required computations. However, the symbolic computations are discrete and not easily parallelizable on modern hardware like GPUs. Second, larger datasets also compound the computational cost of neurosymbolic learning. Deep learning typically addresses this challenge by batching computations across multiple data samples. However, in neurosymbolic learning, the computations may differ across data samples, making it difficult to batch them effectively.
To address these challenges, we need to fundamentally rethink the design of a neurosymbolic framework. One approach is to develop specialized and low-level primitives that scale specific benchmarks but make it time-intensive for developers to write neurosymbolic programs tailored to particular tasks. Alternatively, providing high-level primitives—such as a logic programming language like Scallop (Li et al., 2023) or DeepProbLog (Manhaeve et al., 2018)—simplifies the development of symbolic programs but limits the fine-grained control needed to scale specific applications. Finally, to truly democratize neurosymbolic programming, it is crucial to develop a framework that seamlessly integrates into the everyday deep learning workflows that developers already use.
@@
In this work, we propose Dolphin, a novel framework for scalable neurosymbolic learning. In Dolphin, we build three key components that effectively tackle the scalability and programmability challenges described above. First, we develop a general symbolic representation that efficiently captures the relationships between neural network outputs and associated discrete symbols. Second, we introduce a set of primitives to map forward chaining in symbolic programs to vectorized computations over these representations. Third, we develop a set of vectorized provenance semirings (Green et al., 2007) that are easily pluggable into Dolphin and enable to efficiently compute symbolic gradients. As illustrated in Figure 1(b), these components together allow Dolphin to build a computation graph that spans both symbolic and neural operations, is highly parallelizable, and end-to-end differentiable on GPUs. Finally, Dolphin is implemented as a library that is integrated with PyTorch, allowing users to easily incorporate it into their existing deep learning pipelines.
@@
We evaluate Dolphin on a diverse set of neurosymbolic tasks that involve text, image, video, and multi-modal data, and use rich reasoning features such as recursion and black-box Python functions. Neurosymbolic programs written using Dolphin only require $0.33$%-$37.17$% ($2.77$% on average) of the time to train compared to state-of-the-art baselines including differentiable reasoning frameworks like Scallop, and sampling-based frameworks like ISED and IndeCateR+ while maintaining similar levels of accuracy. We also observe that Dolphin efficiently scales to more complex benchmarks and larger datasets whereas the baselines either time out after 10 hours or fail to converge.
@@
  • We propose Dolphin, a novel neurosymbolic programming framework for end-to-end differentiable symbolic reasoning in a scalable manner.
  • We develop novel abstractions to represent symbolic and neural computations and introduce vectorized primitives for neurosymbolic programs.
  • We develop vectorized provenances that can be plugged into Dolphin for efficient computation of symbolic gradients on parallelizable hardware such as GPUs.
  • We evaluate Dolphin on a diverse range of challenging neurosymbolic tasks across different domains and show that it effectively scales with increasing problem complexity and dataset size.
2 Overview
We illustrate our approach using the MNIST Sum-$N$ task from (De Smet et al., 2024). The goal is to train a model that takes as input $N$ images of MNIST digits and returns the sum of the digits represented by the images. During learning, supervision is provided only on the sum instead of the labels of the digits. The difficulty of the problem scales exponentially as there are $10^{N}$ states in the input space. Further, there are only $19N+$ 1 possible labels, resulting in very sparse supervision.
1class MNISTNet(nn.Module):
2 def __init__(self):
3 super(MNISTNet, self).__init__()
4 ...
5
6class SumNNet(nn.Module):
@@ 2 Overview
12 digits = range(10)
13 D_res = Distribution(self.CNN(imgs[0]), digits)
@@
15 D_i = Distribution(self.CNN(imgs[i]), digits)
@@ 16 D_res = apply(D_res, D_i, lambda x,y: x + y)
17 l_res = get_logits(D_res)
[…]
34
35
36
37
38
39
40
41
42
[…]
49
50In order to support training, the symbolic program must reason over all the outputs of the CNN, and return probability distributions over all the possible results (\$0\$ to \$9N\$).
[…]
52The batched nature of data in machine learning further complicates these calculations since the probabilities of symbols can be different across samples within the same batch.
53As a result, writing neurosymbolic programs in native PyTorch is tedious even for simple tasks.
54
[…]
59
[…]
65This is repeated iteratively until all the outputs of the CNN are summed appropriately.
[…]
67
@@ 3 The Dolphin Framework
3.1 Dolphin Core Design Principles
We based Dolphin’s framework design on the following core principles:
  • Flexible programmability: The framework should allow developers to write neurosymbolic applications in Python with minimal effort, providing intuitive primitives that seamlessly integrate with Python’s rich and expressive language features.
  • End-to-end differentiability on GPUs: The framework should allow any neurosymbolic program to be end-to-end differentiable on GPUs irrespective of the task characteristics.
  • Scalable: The framework should easily scale with greater problem and data complexity.
  • Tunable: Similar to hyperparameters in deep learning, the framework should provide a simple interface for developers to choose provenances (and their configurations) or define new ones.
Next, we show how we realize these principles by describing the key components of Dolphin.
@@ 3.2 The Dolphin Syntax
Dolphin provides a programming interface that developers can use to express symbolic programs in a Pythonic manner. Dolphin maps each operation of the symbolic program to PyTorch which enables end-to-end GPU-accelerated differentiable reasoning. Figure 3 presents Dolphin’s programming interface including the symbolic abstractions and operations over them.
$\begin{array}[]{rcrl}\text{Symbol}&::&s&\in\ \ S\ \ \text{ ({objects})}\\ \text{Tag}&::&t&\in\ \ T\ \ \text{ ({tensors})}\\ \text{Distribution}&::&D&\in\ \ \mathbb{D}=S\rightarrow T\end{array}$
@@
Figure 3: Formal definition of Dolphin’s programming abstractions and primitives.
@@ 3.2.1 Abstractions
The three main abstractions provided by Dolphin for expressing differentiable symbolic programs are shown on the left of Figure 3. Symbols $S$ represent symbolic entities relevant to the program. These entities can be any Pythonic object, such as hand-written digits in MNIST-SumN or coordinates of points in PathFinder. Tags $T$ are tensors that represent relative likelihoods. Typically, tags for symbols are derived from the outputs of machine learning models, such as the probability distribution over digits produced by the CNN classifier in MNIST-SumN. Finally, Distribution $D$ represents the likelihood of an input being classified as one of the pre-defined symbols.
Distributions serve as the fundamental datatype of a Dolphin program and act as its main interface with a machine learning model. For instance, when the developer instantiates a Distribution object, such as in the following code snippet from Figure LABEL:fig:mnist_sum_n_code:
$\texttt{D}_{\text{res}}$ = Distribution(self.CNN(imgs[0]), digits)
the output of the CNN model is directly passed to the Distribution object, effectively acting as an input to the symbolic program. The Distribution object itself, as shown in Figure LABEL:fig:mnist_sum_breakdown, contains batches of tags extracted from the model outputs, and maintains the set of corresponding symbols.
To enable such a seamless integration between the PyTorch model and the symbolic program, Distributions are designed to operate directly over PyTorch tensors. This has two main advantages. First, it preserves the gradients of the model output throughout the symbolic program, enabling end-to-end differentiability. Second, it allows Dolphin to perform operations over an entire batch of tags, leveraging the vectorized operations provided by PyTorch. Dolphin can thus operate efficiently on specialized hardware like GPUs, allowing the symbolic program to scale effectively.
@@ 3.2.2 Operations
Figure 3 shows the five operations supported by Dolphin that developers can use to manipulate Distributions and express complex symbolic programs. We now expand on these operations.
@@
This is the primary operation that developers can use to manipulate Distributions. It takes as inputs $K$ Distributions, where $K\geq 1$, along with a function $f$ of the same arity. This function defines operations over the symbols of $K$ distributions. Apply then computes the results of $f$ over all possible combinations of arguments sourced from the symbols of the Distributions as well as their associated tags, and returns a new Distribution with these results and tags.
This operation occurs in two stages akin to the popular map-reduce pattern. In the map stage, Apply computes the results of $f$ over the symbols of the input Distributions and their associated tags:
$R\ =\ \{\ (f(s_{1},s_{2},\ldots,s_{k}),(t_{1}\otimes t_{2}\otimes\ldots\otimes t_{k}))\ |\ D_{i}(s_{i})=t_{i},i=1,\ldots,k\ \}$ (1)
Here, the tag of each result symbol $f(s_{1},s_{2},\ldots,s_{k})$ is the conjunction $\otimes$ of the tags $(t_{1},t_{2},\ldots,t_{k})$ of the input symbols it was derived from.
Note that the function $f$ can be any user-defined Python function, including complex functions like regex parsing, image processing, or Python’s eval(). As a result, $f$ may be a many-to-one function and the tags shared by a resulting symbol must be aggregated to form the final tags of the output Distribution. We, therefore, shuffle the results from the map stage to compute a function $M$ from each symbol to tags from $R$ associated with it:
@@
Since the tags here are PyTorch tensors representing probabilities, the implementations of the conjunction and disjunction operations are dictated by the underlying provenance used by the program. A more detailed explanation of the provenances is provided in Section 3.3.
Filter. The Filter operation is used to filter out symbols from a Distribution based on some condition. It takes in a single Distribution, along with a user-defined function that returns a boolean value, which acts as the condition. This operation then returns a new Distribution that contains only the symbols that satisfy the condition, along with their tags.
ApplyIf. The ApplyIf operation is a conditional version of Apply. It takes in $K$ Distributions and two functions $f_{\textit{apply}}$ and $f_{\textit{cond}}$ of the same arity. For each combination of symbols from the $K$ Distributions, ApplyIf computes the results of $f_{\textit{apply}}$ and its associated tags only if the condition $f_{\textit{cond}}$ is satisfied over that combination of symbols. The operation then returns a new Distribution with these results and tags.
Union. The Union operation is used to combine two Distributions. It takes in two Distributions and returns a new Distribution that contains the union of the symbols from the two input Distributions, along with their tags. Any symbols common to both input Distributions have their tags merged via a disjunction operation. Union is especially useful when writing recursive programs in Dolphin that require combining the results of multiple recursive calls.
GetProbs. The GetProbs operation is used to extract the probabilities from the tags of a Distribution. This is used mainly once the symbolic program has been executed to extract the final probabilities of the symbols in the output Distribution. These probabilities can then be used to compute the loss function for training the neural components of the neurosymbolic model. The actual extraction of the probabilities from the tags depends on the specific provenance used in the program, which we discuss next.

### v2

@@ Abstract
Neurosymbolic learning enables the integration of symbolic reasoning with deep learning but faces significant challenges in scaling to complex symbolic programs, large datasets, or both. We introduce Dolphin, a framework that tackles these challenges by supporting neurosymbolic programs in Python, executing complex symbolic reasoning on the CPU while vectorizing probabilistic computations and gradient propagation on the GPU. Across 13 benchmarks spanning tasks over text, image, and video data, with symbolic reasoning features like recursion and black-box functions, Dolphin converges to state-of-the-art accuracies on the more complex benchmarks while existing frameworks such as Scallop, ISED, and IndeCateR+ fail to converge within the time limit. On simpler benchmarks, Dolphin matches their performance, while achieving these results 1.71x to 62x faster than the baselines. Overall, Dolphin advances the scalability of neurosymbolic frameworks, achieving state-of-the-art efficiency and convergence on difficult benchmarks where existing frameworks struggle.
@@ 1 Introduction
Figure 1: Illustration of the execution of a neurosymbolic program for the kinship reasoning task CLUTRR. While existing neurosymbolic frameworks run the neural models on the GPU, they run the symbolic program entirely on either the CPU (e.g. Scallop) or the GPU (e.g. Logic Tensor Networks), rendering them inefficient in terms of compute and memory, respectively. In Dolphin, both neural and symbolic components are specified as PyTorch modules, but only probabilistic computations ($\mathcal{P}$) are vectorized on GPU whereas symbolic computations ($\mathcal{L}$) execute on CPU.
@@
Deep learning has made great strides in tasks such as image classification, speech recognition, and natural language processing. With the emergence of foundation models like GPT-4 and SAM, deep learning is increasingly applied to more complex tasks. Despite significant strides, these models remain limited in their ability to reliably perform reasoning required for tasks involving structure, logic, and planning, where symbolic approaches traditionally excel (Kambhampati et al., ). Neurosymbolic programming (Chaudhuri et al., 2021) has emerged as a promising paradigm to incorporate symbolic reasoning into deep learning models, providing the best of both worlds.
@@
Various frameworks have been developed to improve the programmability and accessibility of neurosymbolic applications (Manhaeve et al., 2018; Li et al., 2023; Solko-Breslin et al., 2024). These frameworks support complex symbolic reasoning features like recursion and black-box functions, implement efficient differentiable reasoning algorithms, and provide bindings for deep learning frameworks like PyTorch. However, they incur significant overhead during training.
Figure 1 shows an example of a kinship reasoning task called CLUTRR whose goal is to infer the relationship between two people based on a passage describing interactions and relationships within a family. A natural neurosymbolic formulation splits this task’s computation into a neural component, which extracts relationships from the input passage, and a symbolic component which infers new relationships to obtain the final answer. The latter involves specifying complex manipulations over symbols ($\mathcal{L}$), e.g., multi-hop kinship reasoning between pairs of family members, and performing probabilistic computations ($\mathcal{P}$) to track the probabilities of the symbols derived using $\mathcal{L}$. In general, as the complexity of the symbolic program increases, the number of possible results and their associated weights also grows exponentially, leading to a combinatorial explosion in the number of required computations. This issue is exacerbated by larger datasets usually found in deep learning tasks. Deep learning frameworks typically address this challenge by batching computations across multiple data samples.
Neurosymbolic frameworks like LYRICS (Marra et al., 2019) and Logic Tensor Networks (LTN) (Badreddine et al., 2022) also batch the computations of both $\mathcal{L}$ and $\mathcal{P}$ on the GPU. LTN grounds all logits and discrete symbols as tensors, and the aforementioned computations ($\mathcal{L}$ and $\mathcal{P}$) are specified in differentiable first-order logic as operations over those tensors. These programs output a value quantifying the satisfiability of model outputs with respect to logic constraints. This approach is highly performant for smaller tasks, such as MNIST SumN, where the goal is to predict the sum of $N$ MNIST images. As we see in Table 3 in §$\S$ 3.3, LTN takes around 90 seconds to converge for $N=5$. However, when the complexity increases to $N=10$, LTN runs out of memory on consumer-grade GPUs (here, with a capacity of 11GBs), due to the combinatorial explosion of required symbols (from $10^{5}$ to $10^{10}$) and their probabilities that require to be grounded on the GPU.
On the other hand, neurosymbolic frameworks like DeepProbLog (Manhaeve et al., 2018) and Scallop (Li et al., 2023) run neural models on the GPU but use a separate CPU-based backend for executing both $\mathcal{L}$ and $\mathcal{P}$. This avoids issues of memory consumption on the GPU, but the lack of batched computations on CPU results in slowdowns as the problem complexity increases. We see this in MNIST SumN (Table 2). Scallop requires around 15 minutes to converge for $N=5$, but needs around 1 and 2 hours to converge for $N=10$ and $N=15$, respectively.
@@
In this paper, we propose Dolphin as a solution for scaling neurosymbolic learning. In Dolphin, we build three key components that effectively tackle scalability challenges with existing neurosymbolic frameworks. First, we develop a general symbolic representation that efficiently captures the relationships between neural network outputs as PyTorch tensors on GPU and associated discrete symbols as Python objects on CPU. Second, we introduce a set of primitives to enable writing symbolic manipulations that can be mapped to computations over these representations, while allowing support for black-box Python functions that simplify the writing of complex symbolic programs. Third, we develop a set of vectorized provenance semirings (Green et al., 2007) that are easily pluggable into Dolphin and enable to efficiently compute symbolic gradients.
Together, these components allow Dolphin to build a computation graph that spans both the neural and probabilistic computations ($\mathcal{P}$) that is highly parallelizable, and end-to-end differentiable on GPU. At the same time, it runs $\mathcal{L}$ over discrete symbols on CPU to better manage GPU memory. This allows Dolphin to scale effectively to complex problems such as CLUTRR-N (Table 2) where $N$ denotes the max length of the reasoning chain in the training dataset. In the case of Scallop, as the length of the reasoning chain increases, the gap between convergence times dramatically widens. In contrast, for $N=3$, Dolphin takes around 13 minutes to converge, about 5x faster than Scallop, while for $N=4$, Dolphin takes around 15 minutes, about 8.5x faster than Scallop. Finally, Dolphin is implemented as a library integrated with PyTorch, allowing users to easily incorporate it into their existing deep learning pipelines.
@@
We evaluate Dolphin on a diverse set of neurosymbolic tasks involving text, image, and video, using rich reasoning features like recursion and black-box Python functions. On simpler problems, neurosymbolic programs written using Dolphin match the accuracy of state-of-the-art methods, while achieving these results 47x, 62x, 8x, and 1.7x faster than baselines like Scallop, sampling-based frameworks like ISED and IndeCateR+, and solely GPU based methods like LTN respectively. We also observe that Dolphin efficiently scales to more complex benchmarks and larger datasets, achieving state-of-the-art accuracies. While baselines fail to converge on 5 out of 8 such benchmarks within 10 hours, Dolphin requires 5.5 hours in the worst case.
@@
  • We propose Dolphin, a novel neurosymbolic programming framework for end-to-end differentiable symbolic reasoning in a scalable manner (§2).
  • We develop novel Pythonic abstractions and primitives to enable writing complex symbolic manipulations for neurosymbolic programs (§2.1).
  • We design Dolphin to be extendable to new provenances and develop vectorized provenances that can be plugged into Dolphin for efficient computation of symbolic gradients on parallelizable hardware such as GPU (§2.2).
  • We evaluate Dolphin on a diverse range of 13 challenging neurosymbolic tasks across different domains and show that it effectively scales with increasing problem complexity and dataset size (§3).
1class SumNNet(torch.nn.Module):
@@
7 d = range(10)
8 D_res = Distribution(self.CNN(imgs[0]), d)
@@
10 D_i = Distribution(self.CNN(imgs[i]), d)
@@ 11 D_res = apply(D_res, D_i, lambda x,y: x + y)
42
[…]
49
50To support training, the symbolic program must track digit probabilities, compute probability distributions over all possibilities (\$0\$ to \$9N\$), and propagate gradients for backpropagation. Batched computations further complicate this, making native PyTorch implementations cumbersome.
51
[…]
55
56The programmer can now express the symbolic program in terms of operations over Distributions.
[…]
60This is repeated iteratively until all the outputs of the CNN are summed appropriately.
61
[…]
66
[…]
69This contrasts with systems like Scallop, where tensors are converted into Scallop-friendly tags transferred to a process outside the Python environment with CPU-bound probability computations, restricting scalability.’
@@ 2 The Dolphin Framework
We based Dolphin’s design on four core principles. First is flexible programmability, to allow writing complex symbolic manipulations ($\mathcal{L}$) with Python’s rich and expressive language features. Second, probabilistic computations ($\mathcal{P}$) must allow end-to-end differentiability on the GPU. Third, Dolphin must be scalable to tasks with large data and problem complexity. Finally, it must be tunable, allowing developers to define and choose provenances, treating them like deep learning hyperparameters.
Together, these principles help address the challenges of scaling neurosymbolic frameworks. Flexible programmability and tunability allow us to write complex neurosymbolic programs, while GPU differentiability and scalability work towards tackling problem and data complexity. We now describe Dolphin and show how we realize these principles.
@@ 2.1 The Dolphin Syntax
To allow flexible programmability, Dolphin provides an interface that developers can use to express symbolic programs in a Pythonic manner.
$\begin{array}[]{rcrl}\text{Symbol}&::&s&\in\ \ S\ \ \text{ ({\footnotesize objects})}\\ \text{Tag}&::&t&\in\ \ T\ \ \text{ ({\footnotesize tensors})}\\ \text{Distribution}&::&D&\in\ \ \mathbb{D}=S\rightarrow T\end{array}$
@@
Figure 2: Formal definition of Dolphin’s programming abstractions (top) and primitives (bottom).
@@ 2.1.1 Abstractions
Dolphin provides three main abstractions for expressing symbolic programs, shown in Figure 2. Symbols $S$ represent symbolic entities relevant to the program. These entities can be any Pythonic object, such as hand-written digits in MNIST-SumN or coordinates of points in PathFinder. Tags $T$ are tensors that represent their likelihoods. Typically, tags for symbols are derived from the outputs of machine learning models, such as the logits produced by the CNN classifier in MNIST-SumN. Finally, Distribution $D$ maps a collection of symbols to their corresponding tags.
Distributions serve as the fundamental datatype of a Dolphin program and act as its main interface with a PyTorch model. As seen in the following code snippet from Figure LABEL:fig:mnist_sum_n_code:
$\texttt{D}_{\text{res}}$ = Distribution(self.CNN(imgs[0]), d)
the logits output by the model are directly passed to the Distribution object, effectively acting as an input to the symbolic program. These logits form the batched tags within a Distribution object which also maintains the set of corresponding symbols d. The symbolic manipulations in a Dolphin program occur over the discrete symbols, while the probabilistic computations occur over tags stored as PyTorch tensors. This enables a seamless integration between the PyTorch model and the symbolic program.
This has several advantages. First, it preserves the gradients of the model output throughout the symbolic program, enabling end-to-end differentiability via PyTorch’s autograd, addressing the second core principle of Dolphin. Second, it allows Dolphin to perform operations over an entire batch of tags as per the principle of scalability, leveraging the vectorized operations provided by PyTorch. Dolphin can thus operate efficiently on specialized hardware like GPUs, allowing the symbolic program to scale effectively. Third, since symbol and tag computations are effectively decoupled, operations over symbols can be run on the CPU, allowing the support for arbitrary Python objects and functions in the symbolic program, even while the tag computations are performed on GPU.
@@ 2.1.2 Operations
Dolphin provides five operations to allow the expression of complex neurosymbolic programs in conjunction with user-defined functions, shown in Figure 2.
@@
This is the primary operation that can be used to manipulate Distributions. It takes as inputs $K\geq 1$ Distributions, along with a function $f$ of the same arity. This function defines operations over the symbols of $K$ distributions. Apply then computes the results of $f$ over all possible combinations of arguments sourced from the symbols of the Distributions as well as their associated tags, and returns a new Distribution with these results and tags. This occurs in two stages akin to the popular map-reduce pattern. In the map stage, Apply computes the results of $f$ over the symbols of the input Distributions and conjuncts their tags:
$\begin{split}R\ =\ &\{\ (f(s_{1},s_{2},\ldots,s_{k}),(t_{1}\otimes t_{2}\otimes\ldots\otimes t_{k}))\ |\\ &D_{i}(s_{i})=t_{i},i=1,\ldots,k\ \}\end{split}$ (1)
Here, the tag of each result symbol $f(s_{1},s_{2},\ldots,s_{k})$ is the conjunction $\otimes$ of the tags $(t_{1},t_{2},\ldots,t_{k})$ of the input symbols it was derived from. The function $f$ is executed sequentially on the CPU for each combination of symbols as function $f$ can be any user-defined Python function, including complex control flows and operations like regex parsing, image processing, or Python’s eval(). It may also be a many-to-one function and the tags shared by a resulting symbol must be aggregated to form the final tags of the output Distribution. We, therefore, shuffle the results from the map stage to compute a function $M$ from each symbol to tags from $R$ associated with it:
@@
Since the tags here are PyTorch tensors representing probabilities, the implementations of the conjunction and disjunction operations are dictated by the underlying provenance specified by the program. A more detailed explanation of the provenances is provided in Section 2.2.
Filter. The Filter operation is used to filter out symbols from a Distribution. It takes in a single Distribution, along with a user-defined function that returns a boolean value. This operation then returns a new Distribution that contains only symbols that satisfy the condition with their tags.
ApplyIf. This operation is a conditional version of Apply. It takes in $K$ Distributions and functions $f_{\textit{apply}}$ and $f_{\textit{cond}}$ of the same arity. For each combination of symbols from the $K$ Distributions, ApplyIf computes $f_{\textit{apply}}$ and its associated tags only if the condition $f_{\textit{cond}}$ is satisfied over that combination of symbols. The operation then returns a new Distribution with these results and tags.
Union. The Union operation takes in two Distributions and returns a new Distribution containing the union of the input symbols, along with their tags. Any symbols common to both input Distributions have their tags disjuncted.
GetProbs. The GetProbs operation extracts the probabilities from the tags of a Distribution. This is used mainly once the symbolic program has been executed to extract the final probabilities of the symbols in the output Distribution. These probabilities can then be used to compute the loss function for training the neural model.
2.1.3 Writing Complex Dolphin Programs
Some neurosymbolic tasks require the writing of programs containing complex control flows and recursion (e.g. the PathFinder task, §$\S$ 3.1). This can be done in one of two ways. The simplest way is to specify any control flow operations within the user-defined functions supplied to the Dolphin operations. Alternatively, one can specify branches of control flow separately and merge their results via Union, as shown in Figure 4 (Appendix B for more details).
In some cases, even though Dolphin limits the effect of combinatorial explosion in terms of efficiency and memory usage, the number of combinations may still be excessive. In such cases, Dolphin allows developers to sample subsets of symbols from Distribution objects specified within the symbolic program, effectively limiting the number of symbols processed in each operation.